# Distilling the Knowledge in Data Pruning

### v1

@@ Abstract
Keywords:
Data pruning Knowledge distillation Image classification
@@ 1 Introduction
(a) In knowledge distillation for model compression (left), a large teacher network is used to guide the training of a smaller student network. In contrast, here we investigate the usage of a teacher model, pre-trained on a full dataset, to guide a student model during training on a pruned subset of the data (right).
@@
Figure 1: Knowledge distillation for data pruning. (a) The difference between KD for model compression and KD for data pruning. (b) We find that by integrating KD into the training, simple random pruning outperforms other sophisticated pruning algorithms across all pruning regimes. (c) Interestingly, we observe that when using small data fractions, training with large teachers degrades accuracy, while smaller teachers are favored. This suggests that in high pruning regimes (low $f$), the training is more sensitive to the capacity gap between the teacher and the student.
@@
In this paper, we explore the benefit of using a model trained on a complete dataset to enhance training on a pruned subset of the data using knowledge distillation (KD). The motivation behind this approach is that a teacher model trained on the complete dataset captures essential information and core statistics about the entire data. This knowledge can then be utilized when training on a pruned subset. While KD has been extensively studied and demonstrated significant improvements in tasks such as model compression, herein we aim to investigate its impact in the context of data pruning and propose innovative findings for practical usage. Note that, in contrast to traditional model compression techniques, our approach focuses on self-distillation (SD), where the teacher and student have identical architectures. An overview of the proposed training scheme is presented in Fig. 0(a).
@@
  • We explore the utilization of knowledge distillation while training on pruned data by leveraging the soft predictions of a teacher pre-trained on the complete data. Using KD, we demonstrate consistent improvement across all pruning regimes, various pruning algorithms, and on multiple datasets.
@@
  • We find that KD is robust to the choice of the pruning mechanism for high pruning fractions. In particular, using KD, random pruning achieves comparable or superior accuracy compared to other sophisticated methods across all pruning regimes.
@@ 2 Related work
Data pruning. Data pruning, also known as coreset selection [27, 18, 37], refers to methods aiming to reduce the dataset size for training neural networks. Recent approaches have shown significant progress in retaining less data while maintaining high classification accuracy [38, 32, 10, 25, 5, 35]. In [35], the authors showed theoretically and empirically that data pruning can improve the power law scaling of the dataset size by choosing an optimal pruning fraction as a function of the initial dataset size. Additionally, studies in [35, 2] have demonstrated that existing pruning algorithms often underperform when compared to random pruning methods, especially in high pruning regimes. Recently, in [48], the authors suggested a theoretical explanation to this accuracy drop, and proposed a coverage-centric pruning approach which better handles the data coverage.
Data pruning proves valuable at reducing memory and computational cost in various applications, including tasks such as hyper-parameter search [6], neural architecture search [8], continual and incremental learning [22], as well as active learning [27, 5].
@@
In our paper, we explore the process of distilling knowledge from a model trained on a large dataset to a model trained on a pruned subset of the original data. Specifically, we focus on self-distillation and present several striking observations that emerge when integrating SD for data pruning.
@@ 3.1 Training on the pruned dataset using KD
Integrating the KD loss into the training process allows us to leverage the valuable knowledge embedded in the teacher’s soft predictions $q_{i}$. These predictions may encapsulate potential relationships between categories and class hierarchies, accumulated by the teacher during its training on the entire dataset. To illustrate this, we provide a qualitative example in Fig. 2 that presents the soft predictions generated for a specific sample from the CIFAR-100 dataset. CIFAR-100 comprises 100 classes, organized into 20 super-classes, each containing 5 sub-classes. For example, the super-class "People" contains the classes: "Baby", "Boy", "Girl", "Man", and "Woman". As shown in Fig. 2 (top), the teacher accurately predicts the ground-truth class "Girl" (class index 35) with high confidence while also assigning high confidence values to the classes "Woman" (98), "Man" (46), and "Boy" (11). This ‘dark knowledge’ is valuable for training as it offers a broader view of class hierarchies and data distribution. Fig. 2 (middle) illustrates that a model trained on only $25\%$ of the data fails to capture such class relationships. Intuitively, reliable data and class distributions can be effectively learned from large datasets, but are harder to infer from small datasets. Conversely, in Fig. 2 (bottom) we show that using knowledge distillation, the student successfully learns these delicate data relationships from the teacher despite training only on the pruned data.
Figure 2: Learning from the teacher predictions. An example of soft predictions computed by a teacher model trained on the entire data (top), a model trained on $25\%$ of the data (middle), and a student model trained on $25\%$ of the data with KD (bottom), for an evaluation sample of class "Girl" from CIFAR-100. Using KD, the student can better learn close or ambiguous categories by leveraging knowledge captured by the teacher from the full dataset.
@@ 3.2 Mitigating noisy samples in pruned datasets
Refer to caption (a) CIFAR-100 highest score pruning samples Refer to caption (b) SVHN highest score pruning samples
Figure 3: Highest scoring samples. Top 10 highest scoring samples selected by the ‘forgetting’ method for CIFAR-100 and SVHN datasets. The labels of the majority of the images are ambiguous or hard to resolve due to class complexity or low image quality.
@@ 4 Experimental results
Implementation Details. For computational efficiency we conduct our self-distillation experiments on all datasets using the ResNet-32 [13] architecture, except for ImageNet for which we utilize the larger ResNet-50. Our training and distillation recipes are simple. We utilize SGD with Momentum to optimize the models and incorporate basic data-augmentations during training. Additional implementation details can be found in the supplementary.
@@ 4.1 Training on pruned data with KD
As can be observed in Figs. 4 and 5, the incorporation of KD into the training process consistently enhances model accuracy across all of the tested scenarios, regardless of the tested dataset, pruning method or pruning level. For example, compared to baseline models trained on the full datasets without KD, utilizing KD can lead to comparable accuracy levels by retaining only small portions of the original datasets (e.g., 10%, 30%, 50% on SVHN, CIFAR-10, and CIFAR-100, respectively, using ‘forgetting’). In fact, even on a large scale dataset as ImageNet, comparable accuracy can be achieved by randomly retaining just 30% of the data, while training on larger subsets remarkably results in superior accuracy to the baseline (e.g., +1.6% using a random subset of 70%). In contrast, we observe that in the absence of KD, preserving accuracy is only possible in the presence of more data (i.e., more relaxed pruning levels). In particular, on CIFAR-100 we found accuracy preservation to be unattainable without KD even in the least aggressive pruning scenario we have tested ($f=0.9$), and regardless of the pruning method.
@@
Figure 7: Exploring the effect of the teacher’s capacity. Accuracy results for a student with (a) ResNet-32 and (b) ResNet-20 architectures while using teacher models with increasing capacities along the horizontal axes. In each instance, we denote the teacher whose architecture matches that of the student by ‘SD’ (self-distillation). We use random pruning with different fractions. Surprisingly, under low pruning factors, increasing the teacher’s capacity results in lower student accuracy.
@@ 4.3 Using teachers of different capacities
Until now, we have focused on the case where both the student and teacher share the same architecture (i.e., self-distillation). In this section, we explore how the capacity of the teacher affects the student’s performance across different pruning regimes. In Fig. 6(a), we present accuracy results across various pruning factors for the case of randomly pruning CIFAR-100 and training with a ResNet-32 student. We employ 6 teacher architectures of increasing capacities: (1) ResNet-14 with $69.9\%$ accuracy, (2) ResNet-20 with $70.23\%$ accuracy, (3) ResNet-32 with $71.6\%$ accuracy, (4) ResNet-56 with $72.7\%$ accuracy, (5) ResNet-110 with $74.4\%$ accuracy, and (6) WRN-40-2 with $75.9\%$ accuracy. Also, note that for each teacher architecture we experiment with five different temperature values in the range $2-7$. We show the impact of the temperature selection in the supplementary. Similarly, in Fig. 6(b) we present results for the same experiment using a ResNet-20 student, while Fig. 6(c) depicts results of a similar experiment on CIFAR-10 for the ResNet-32 student.
As observed, at low pruning factors, increasing the teacher’s capacity harms the accuracy of the student. Conversely, at high pruning factors (e.g. $0.9$ and $1$) larger teachers generally lead to improved results, as expected in common KD for model compression. This trend is consistently observed across various student architectures and datasets, and is robust to the selection of the KD temperature. Additional results are provided in the supplementary.
This observation highlights a striking phenomenon: the capacity gap problem, which denotes the disparity in architecture size between the teacher and student, becomes more pronounced when applying knowledge distillation during training on pruned data. Surprisingly, for small pruning fractions, teachers with a smaller capacity than the student can lead to higher student accuracy levels compared to larger teachers. The results also suggest that the effectiveness of smaller teachers in improving the student is bounded to an extent by their own accuracy levels.
@@ 5 Conclusion
We believe that these empirical observations can pave the way for in-depth research, both theoretical and experimental, to enhance our understanding of the optimization process when training models on pruned data using KD.

### v2

@@ 1 Introduction
(a) KD for data pruning
@@
Figure 1: Knowledge distillation for data pruning. (a) We investigate the usage of a teacher model, pre-trained on a full dataset, to guide a student model during training on a pruned subset of the same data. (b) We find that by integrating KD into the training, simple random pruning outperforms other sophisticated pruning algorithms across all pruning regimes. (c) Interestingly, we observe that when using small data fractions, training with large teachers degrades accuracy, while smaller teachers are favored. This suggests that in high pruning regimes (low $f$), the training is more sensitive to the capacity gap between the teacher and the student.
@@
In this paper, we explore the benefit of using a model trained on a complete dataset to enhance training on a pruned subset of the data using knowledge distillation (KD). The motivation behind this approach is that a teacher model trained on the complete dataset captures essential information and core statistics about the entire data. This knowledge can then be utilized when training on a pruned subset. While KD has been extensively studied and demonstrated significant improvements in tasks such as model compression, herein we aim to investigate its impact in the context of data pruning and propose innovative findings for practical usage. Note that, in contrast to traditional model compression techniques, here we focus on self-distillation (SD), where the teacher and student have identical architectures. The training scheme is illustrated in Fig. 1(a).
@@
  • Utilizing KD in data pruning, we find that training is robust to the choice of pruning mechanism at high pruning fractions. Notably, random pruning with KD achieves comparable or superior accuracy compared to other sophisticated methods across all pruning regimes.
@@ 2 Related work
Data pruning. Data pruning, also known as coreset selection [27, 18, 37], refers to methods aiming to reduce the dataset size for training neural networks. Recent approaches have shown significant progress in retaining less data while maintaining high classification accuracy [38, 32, 10, 25, 5, 35]. In [35], the authors showed theoretically and empirically that data pruning can improve the power law scaling of the dataset size by choosing an optimal pruning fraction as a function of the initial dataset size. Additionally, studies in [35, 2] have demonstrated that existing pruning algorithms often underperform when compared to random pruning methods, especially in high pruning regimes. Recently, in [49], the authors suggested a theoretical explanation to this accuracy drop, and proposed a coverage-centric pruning approach which better handles the data coverage. Also, in [41], the authors proposed to model the sample selection procedure as a constrained discrete optimization problem.
Data pruning proves valuable at reducing memory and computational cost in various applications, including tasks such as hyper-parameter search [6], NAS [8], continual and incremental learning [22], as well as active learning [27, 5].
@@
In our paper, we explore the process of distilling knowledge from a model trained on a large dataset to a model trained on a pruned subset of the original data. We focus on self-distillation and present several striking observations that emerge when integrating SD for data pruning.
@@ 3.1 Training on the pruned dataset using KD
Integrating the KD loss into the training process allows us to leverage the valuable knowledge embedded in the teacher’s soft predictions $q_{i}$. These predictions may encapsulate potential relationships between categories and class hierarchies, accumulated by the teacher during its training on the entire dataset. To illustrate this, we provide a qualitative example in Fig. 2 that presents the soft predictions generated for a specific sample from the CIFAR-100 dataset. CIFAR-100 comprises 100 classes, organized into 20 super-classes, each containing 5 sub-classes. For example, the super-class ”People” contains the classes: ”Baby”, ”Boy”, ”Girl”, ”Man”, and ”Woman”. As shown in Fig. 2 (top), the teacher accurately predicts the ground-truth class ”Girl” (class index 35) with high confidence while also assigning high confidence values to the classes ”Woman” (98), ”Man” (46), and ”Boy” (11). This ‘dark knowledge’ is valuable for training as it offers a broader view of class hierarchies and data distribution. Fig. 2 (middle) illustrates that a model trained on only $25\%$ of the data fails to capture such class relationships. Intuitively, reliable data and class distributions can be effectively learned from large datasets, but are harder to infer from small datasets. Conversely, in Fig. 2 (bottom) we show that using knowledge distillation, the student successfully learns these delicate data relationships from the teacher despite training only on the pruned data.
Figure 2: Learning from the teacher predictions. An example of soft predictions computed by a teacher model trained on the entire data (top), a model trained on $25\%$ of the data (middle), and a student model trained on $25\%$ of the data with KD (bottom), for an evaluation sample of class ”Girl” from CIFAR-100. Using KD, the student can better learn close or ambiguous categories by leveraging knowledge captured by the teacher from the full dataset.
@@ 3.2 Mitigating noisy samples in pruned datasets
(a) CIFAR-100 highest score pruning samples
(b) SVHN highest score pruning samples
Figure 3: Highest scoring samples. Top 10 highest scoring samples selected by the ‘forgetting’ pruning method for CIFAR-100 and SVHN datasets. The labels of the majority of the images are ambiguous due to class complexity or low image quality.
(a) CIFAR-100
(b) SVHN
(c) CIFAR-10
@@ 4.1 Training on pruned data with KD
As can be observed in Figs. 4 and 5, the incorporation of KD into the training process consistently enhances model accuracy across all of the tested scenarios, regardless of the tested dataset, pruning method or pruning level. For example, compared to baseline models trained on the full datasets without KD, utilizing KD can lead to comparable accuracy levels by retaining only small portions of the original datasets (e.g., 10%, 30%, 50% on SVHN, CIFAR-10, and CIFAR-100, respectively, using ‘forgetting’). In fact, even on a large scale dataset as ImageNet, comparable accuracy can be achieved by randomly retaining just 30% of the data, while training on larger subsets remarkably results in superior accuracy to the baseline (e.g., +1.6% using a random subset of 70%).
@@
Figure 7: Exploring the effect of the teacher’s capacity. Accuracy results for a student with (a) ResNet-32 and (b) ResNet-20 architectures while using teacher models with increasing capacities along the horizontal axes. In each instance, we denote the teacher whose architecture matches that of the student by ‘SD’ (self-distillation). We use random pruning with different fractions. Interestingly, under low pruning factors, increasing the teacher’s capacity results in lower student accuracy.
@@ 4.3 Using teachers of different capacities
Until now, we have focused on the case where both the student and teacher share the same architecture (i.e., self-distillation). In this section, we explore how the capacity of the teacher affects the student’s performance across different pruning regimes. In Fig. 7(a), we present accuracy results across various pruning factors for the case of randomly pruning CIFAR-100 and training with a ResNet-32 student. We employ 6 teacher architectures of increasing capacities: (1) ResNet-14 with $69.9\%$ accuracy, (2) ResNet-20 with $70.23\%$ accuracy, (3) ResNet-32 with $71.6\%$ accuracy, (4) ResNet-56 with $72.7\%$ accuracy, (5) ResNet-110 with $74.4\%$ accuracy, and (6) WRN-40-2 with $75.9\%$ accuracy. Also, note that for each teacher architecture we experiment with five different temperature values in the range $2-7$. We show the impact of the temperature selection in the supplementary. Similarly, in Fig. 7(b) we present results for the same experiment using a ResNet-20 student, while Fig. 7(c) depicts results of a similar experiment on CIFAR-10 for the ResNet-32 student. As observed, at low pruning factors, increasing the teacher’s capacity harms the accuracy of the student. This trend is consistently observed across various student architectures and datasets, and is robust to the selection of the KD temperature. Additional results are provided in the supplementary.
This observation highlights a striking phenomenon: the capacity gap problem, which denotes the disparity in architecture size between the teacher and student, becomes more pronounced when applying knowledge distillation during training on pruned data.